# On the multiplier spectrum of polynomials

### v1

@@ 3.1. Pre-simple polynomials and the case $d\geq 4$
By Theorem 3.14, Lemma 3.15, and Lemma 3.16, there exists a non-empty Zariski open subset $W_{d}$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for every $f\in W_{d}$, the polynomial $f$ is pre-simple and
(3.1) $\{g\in{\rm Rat}_{d}(\mathbb{C}):\mu_{g}=\mu_{f}\}=\{f\}.$
We show that $W_{d}$ also satisfies (2). Let $f\in W_{d}$ and $g\in{\rm Poly}^{d}(\mathbb{C})$ be a pre-simple polynomial intertwined with $f$. By Theorem 3.22, there exist $k\in\mathbb{Z}_{>0}$ and $\alpha\in{\rm PGL}_{2}(\mathbb{C})$ such that
@@
By (3.1), $f=\alpha\circ g\circ\alpha^{-1}$. Since $\alpha$ is a degree-$1$ polynomial, $[f]=[g]$ in ${\rm MPoly}^{d}(\mathbb{C})$.
@@ 3.2. The case of degree $2\leq d\leq 3$
Having analyzed general polynomials of degree $d\geq 4$ in §3.1 using pre-simplicity, we now consider general polynomials of degree $2\leq d\leq 3$, largely motivated by Pakovich’s work [Pak25b] on general rational maps of degree $2\leq d\leq 3$. We show that for $d\in\{2,3\}$, there exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for all $(f,g)\in U\times U$, the polynomials $f$ and $g$ are intertwined if and only if they are conjugate. This result is also a direct corollary of [FG22, Theorems 3.51 and 3.52] (see [GNY19, Thoerem 1.4]).
For $d=2$, the moduli space ${\rm MPoly}^{2}\cong\mathbb{A}^{1}$ has dimension $1$, and every $f\in{\rm Poly}^{2}(\mathbb{C})$ is conjugate to a unique polynomial of the form $\phi_{c}(z)=z^{2}+c$ with $c\in\mathbb{C}$. Thus, degree $2$ polynomials are easier to study due to the simple normal forms $\phi_{c}$. For $d=3$, Theorem 3.7 holds, which concerns decompositions of iterates of a general polynomial and relies on Ritt’s theory. For both $d=2$ and $d=3$, we show that such a Zariski open subset $U$ exists.
For a non-constant rational map $f\in\mathbb{C}(z)$, define the groups
$\Sigma(f):=\{\sigma\in{\rm PGL}_{2}(\mathbb{C}):f\circ\sigma=f\}\quad\text{and}\quad\Sigma_{\infty}(f):=\bigcup_{k\geq 1}\Sigma(f^{\circ k}).$
Clearly, $\Sigma(f)\subseteq\Sigma_{\infty}(f)$, and this inclusion is preserved under conjugation: for $\tau\in{\rm PGL}_{2}(\mathbb{C})$,
$\Sigma(\tau\circ f\circ\tau^{-1})=\tau\circ\Sigma(f)\circ\tau^{-1}\subseteq\tau\circ\Sigma_{\infty}(f)\circ\tau^{-1}=\Sigma_{\infty}(\tau\circ f\circ\tau^{-1}).$
@@
For every $c\in\mathbb{C}^{\times}$, we have $\Sigma(\phi_{c})=\Sigma_{\infty}(\phi_{c})=\{z,-z\}$, the cyclic group of order two generated by z-z.
@@
Clearly, $\{z,-z\}\subseteq\Sigma(\phi_{c})\subseteq\Sigma_{\infty}(\phi_{c})$.
Since $c\neq 0$, the map $\phi_{c}$ has a unique totally ramified fixed point at $\infty$. Thus, every $\sigma\in\Sigma(\phi_{c})$ fixes $\infty$, so $\sigma$ is a linear polynomial. Write $\sigma(z)=az+b$ with $a\in\mathbb{C}^{\times}$, $b\in\mathbb{C}$. The equality $\phi_{c}\circ\sigma=\phi_{c}$ implies $\sigma(z)=z$ or $\sigma(z)=-z$. Hence, $\{z,-z\}=\Sigma(\phi_{c})$.
We now show by induction on $n\in\mathbb{Z}_{>0}$ that $\Sigma(\phi_{c})=\Sigma(\phi_{c}^{\circ n})$ for all $n$, so $\Sigma(\phi_{c})=\Sigma_{\infty}(\phi_{c})$. The case $n=1$ is trivial. Assume $n\geq 2$ and $\Sigma(\phi_{c})=\Sigma(\phi_{c}^{\circ(n-1)})$. Clearly, $\Sigma(\phi_{c})\subseteq\Sigma(\phi_{c}^{\circ n})$. To show $\Sigma(\phi_{c}^{\circ n})\subseteq\Sigma(\phi_{c})$, let $\sigma\in\Sigma(\phi_{c}^{\circ n})$. Since $c\neq 0$, $\phi_{c}^{\circ n}$ has a unique totally ramified fixed point at $\infty$, so $\sigma$ fixes $\infty$ and is a linear polynomial: $\sigma(z)=az+b$ with $a\in\mathbb{C}^{\times}$, $b\in\mathbb{C}$. By [Pak25b, Lemma 3.4] and $\phi_{c}\circ\phi_{c}^{\circ(n-1)}=\phi_{c}\circ(\phi_{c}^{\circ(n-1)}\circ\sigma)$, we have either $\phi_{c}^{\circ(n-1)}=\phi_{c}^{\circ(n-1)}\circ\sigma$ or $\phi_{c}^{\circ(n-1)}=-\phi_{c}^{\circ(n-1)}\circ\sigma$. If $\phi_{c}^{\circ(n-1)}=\phi_{c}^{\circ(n-1)}\circ\sigma$, then by induction, $\sigma\in\Sigma(\phi_{c}^{\circ(n-1)})=\Sigma(\phi_{c})$, so $\Sigma(\phi_{c}^{\circ n})\subseteq\Sigma(\phi_{c})$.
Suppose $\phi_{c}^{\circ(n-1)}=-\phi_{c}^{\circ(n-1)}\circ\sigma$. Comparing coefficients:
  • The coefficient of $z^{2^{n-1}}$ gives $a^{2^{n-1}}=-1$.
  • The coefficient of $z^{2^{n-1}-1}$ gives $b=0$.
  • The coefficient of $z^{2^{n-1}-2}$ gives $2^{n-2}c=-a^{2^{n-1}-2}2^{n-2}c$, so $a^{2^{n-1}-2}=-1$ (since $c\neq 0$).
Then
@@
which is a contradiction. Hence, $\Sigma(\phi_{c}^{\circ n})\subseteq\Sigma(\phi_{c})$. By induction, we obtain $\Sigma(\phi_{c})=\Sigma_{\infty}(\phi_{c})$. ∎
@@
There exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{2}(\mathbb{C})$ such that for every $F\in U$:
@@
    $\Sigma(F)=\Sigma_{\infty}(F)$ is cyclic of order $2$;
@@
    if $G(z)\in\mathbb{C}(z)$ is a rational map of degree $2$ such that $G^{\circ k}=F^{\circ k}$ for some $k\in\mathbb{Z}_{>0}$, then $G=F$;
  3. (3)
    for every $l\in\mathbb{Z}_{>0}$, every decomposition of $F^{\circ l}$ into indecomposable rational maps (of degree $\geq 2$) is equivalent to $F^{\circ l}=F\circ\cdots\circ F$.
@@
Let $U:=\{a_{1}^{2}-2a_{1}\neq 4a_{0}a_{2}\}$, a non-empty Zariski open subset of
${\rm Poly}^{2}(\mathbb{C})=\{a_{0}z^{2}+a_{1}z+a_{2}\}\cong\{(a_{0},a_{1},a_{2}):a_{0}\neq 0\}.$
A simple computation shows that $a_{0}z^{2}+a_{1}z+a_{2}\in{\rm Poly}^{2}(\mathbb{C})$ is conjugate to some $z^{2}+c$ with $c\neq 0$ if and only if it lies in $U$.
Let $F(z)=a_{0}z^{2}+a_{1}z+a_{2}\in U$, and let $c=c(F)\in\mathbb{C}^{\times}$ be such that $F$ is conjugate to $\phi_{c}$. Choose $\tau\in{\rm PGL}_{2}(\mathbb{C})$ with $\tau\circ F\circ\tau^{-1}=\phi_{c}$.
(1) By Lemma 3.24, $\Sigma(\phi_{c})=\Sigma_{\infty}(\phi_{c})=\{z,-z\}$ is cyclic of order $2$, so $\Sigma(F)=\Sigma_{\infty}(F)$ is also cyclic of order $2$.
[…]
$\phi_{c}^{\circ k}\circ\mu_{\tilde{G}}=\tilde{G}^{\circ k}\circ\mu_{\tilde{G}}=\tilde{G}^{\circ k}=\phi_{c}^{\circ k},$
so $\mu_{\tilde{G}}\in\Sigma_{\infty}(\phi_{c})=\{\pm z\}$ by Lemma 3.24. Since $\mu_{\tilde{G}}(z)\neq z$, we have $\mu_{\tilde{G}}(z)=-z$. By [Pak25b, Lemma 3.5], $\tilde{G}=\nu\circ\phi_{c}$ for some $\nu\in{\rm PGL}_{2}$. From $\phi_{c}^{\circ k}=\tilde{G}^{\circ k}=(\nu\circ\phi_{c})^{\circ k}$, we get $\phi_{c}^{\circ(k-1)}=(\nu\circ\phi_{c})^{\circ(k-1)}\circ\nu$, so
@@
The same argument shows that the generator of $\Sigma(\phi_{c}\circ\nu)$ is $\mu_{\phi_{c}\circ\nu}(z)=-z$. Note that
$\phi_{c}\circ\nu\circ(\nu^{-1}\circ(-z)\circ\nu)=\phi_{c}\circ(-z)\circ\nu=\phi_{c}\circ\nu,$
so $\nu^{-1}\circ(-z)\circ\nu(z)=\mu_{\phi_{c}\circ\nu}(z)=-z$, hence $\nu(-z)=-\nu(z)$. Thus, $\nu(z)=\lambda z$ or $\nu(z)=\lambda/z$ for some $\lambda\in\mathbb{C}^{\times}$. If $\nu(z)=\lambda/z$, then $\infty$ is not a totally ramified fixed point of $(\lambda/(z^{2}+c))^{\circ k}=\phi_{c}^{\circ k}$ (since $c\neq 0$), contradicting that $\phi_{c}^{\circ k}$ is a polynomial. So $\nu(z)=\lambda z$. Then
@@
Comparing the coefficient of $z^{2^{k}}$ of the two sides of (3.2) gives $\lambda^{2^{k}-1}=1$. If $k=1$, then $\lambda=1$. If $k\geq 2$, the coefficient of $z^{2^{k}-4}$ gives
@@
so $\lambda^{2}=1$ (since $c\neq 0$ and $\lambda^{2^{k}-1}=1$). Then $\lambda=(\lambda^{2})^{2^{k-1}}/\lambda^{2^{k}-1}=1$. Hence, $\nu(z)=z$, so $\tilde{G}=\phi_{c}$ and $G=F$.
(3) We proceed by induction on $l$. For $l=1$, the result is trivial since $\deg(F^{\circ l})=2$ is prime. Assume $l\geq 2$ and the result holds for l1l-1. Let $F^{\circ l}=F_{r}\circ\cdots\circ F_{1}$ be a decomposition into indecomposable rational maps. It suffices to show $F_{1}=\mu\circ F$ for some $\mu\in{\rm PGL}_{2}(\mathbb{C})$. By [Pak25b, Corollary 2.2], $\deg(F_{1})=2$. Let $\mu_{F_{1}}$ be the generator of $\Sigma(F_{1})$ (cyclic of order $2$ by [Pak25b, Lemma 3.4 (i)]). Then
$F^{\circ l}\circ\mu_{F_{1}}=F_{r}\circ\cdots\circ F_{1}\circ\mu=F_{r}\circ\cdots\circ F_{1}=F^{\circ l},$
so ${\rm Id}\neq\mu_{F_{1}}\in\Sigma_{\infty}(F)=\Sigma(F)$. By [Pak25b, Lemma 3.5], $F_{1}=\mu\circ F$. The result follows by induction. ∎
@@
There exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{3}(\mathbb{C})$ such that for every $F\in U$:
@@
    if $G(z)\in\mathbb{C}(z)$ is a rational map of degree $3$ such that $G^{\circ k}=F^{\circ k}$ for some $k\in\mathbb{Z}_{>0}$, then $G=F$;
  3. (3)
    for every $l\in\mathbb{Z}_{>0}$, every decomposition of $F^{\circ l}$ into indecomposable rational maps (of degree $\geq 2$) is equivalent to $F^{\circ l}=F\circ\cdots\circ F$.
@@
By Lemmas 3.15 and 3.16, there exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{3}(\mathbb{C})$ such that every $F\in U$ is pre-simple and $G(F)=1$. Then (3) follows from Theorem 3.7.
Let $F\in U$. We prove (1) and (2).
(1) $F$ is pre-simple by construction. After conjugation, assume $F(z)=z^{3}+az+b$ with $a,b\in\mathbb{C}$. Since $F$ is pre-simple, $F^{\prime}(z)=3z^{3}+a$ has no multiple roots, so $a\neq 0$. Let $n\geq 1$ and $\sigma\in\Sigma(\phi_{c}^{\circ n})$. Then $\sigma$ is a linear polynomial (as in the proof of Lemma 3.24). Write $\sigma(z)=ez+f$ with $e\in\mathbb{C}^{\times}$, $f\in\mathbb{C}$. By induction, the leading terms of $(z^{3}+az+b)^{\circ k}$ are:
(3.3) $(z^{3}+az+b)^{\circ k}=z^{3^{k}}+0\cdot z^{3^{k}-1}+3^{k-1}az^{3^{k}-2}+O(z^{3^{k}-3}).$
Substituting into $(z^{3}+az+b)^{\circ k}\circ(ez+f)=(z^{3}+az+b)^{\circ k}$ and comparing coefficients of $z^{3^{k}},z^{3^{k}-1},z^{3^{k}-2}$ shows $(e,f)=(1,0)$, so $\sigma(z)=z$. Hence, $\Sigma(F)=\Sigma_{\infty}(F)=1$.
(2) Let $G\in\mathbb{C}(z)$ be a rational map of degree $3$ with $G^{\circ k}=F^{\circ k}$ for some $k\in\mathbb{Z}_{>0}$. Since $\deg(F)=\deg(G)=3$ is prime, both are indecomposable. By (3), the decomposition $F^{\circ k}=G\circ\cdots\circ G$ implies $G=\nu\circ F$ for some $\nu\in{\rm PGL}_{2}(\mathbb{C})$. Then $F^{\circ k}=(\nu\circ F)^{\circ k}$, so $F^{\circ(k-1)}=(\nu\circ F)^{\circ(k-1)}\circ\nu$. Composing with $F$ gives $F^{\circ k}=(F\circ\nu)^{\circ k}$. Applying (3) again, $F\circ\nu=\delta\circ F$ for some $\delta\in{\rm PGL}_{2}(\mathbb{C})$. Then $\nu\in G(F)=1$, so $\nu(z)=z$ and $G=F$. ∎
@@
For $d\geq 4$, the result follows from Lemmas 3.18 and 3.15.
If $d=2$, then $F$ is a generalized Lattès map if and only if it is exceptional (Proposition 2.4). Exceptional polynomials of degree $2$ are conjugate to $z^{2}$ or $T_{2}(z)=z^{2}-2$. Since ${\rm MPoly}^{2}(\mathbb{C})\cong\mathbb{C}$ has dimension $1>0$, a general $F$ is not exceptional.
If $d=3$, then $F$ is exceptional if and only if it is conjugate to $z^{3}$ or $\pm T_{3}(z)=\pm(z^{3}-3z)$. Thus a general $F$ is not exceptional. By Proposition 2.4 and $\deg(F)=3$, a non-exceptional $F$ is a generalized Lattès map if and only if it is conjugate to $zR(z)^{2}=z(az+b)^{2}$ for some $a,b\in\mathbb{C}^{\times}$ (since $F$ is not conjugate to $z^{3}$). Thus, $F$ is a generalized Lattès map if and only if there exist $a,b,e\in\mathbb{C}^{\times}$ and $f\in\mathbb{C}$ such that
(3.4) $F(z)=\frac{(ez+f)(a(ez+f)+b)^{2}-f}{e}.$
@@
There are only $3$ parameters $\lambda,b\in\mathbb{C}^{\times}$ and $\mu\in\mathbb{C}$, while ${\rm Poly}^{3}(\mathbb{C})=\{a_{0}z^{3}+a_{1}z^{2}+a_{2}z+a_{3}:(a_{0},a_{1},a_{2},a_{3})\in\mathbb{C}^{\times}\times\mathbb{C}^{3}\}$ has dimension $4>3$. Hence, a general $F$ is not a generalized Lattès map. ∎
@@
Let $d\in\{2,3\}$. There exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for every $F\in U$, if $B,X\in\mathbb{C}(z)\setminus\mathbb{C}$ and $k\in\mathbb{Z}_{>0}$ satisfy $F^{\circ k}\circ X=X\circ B$, then there exist $m\in\mathbb{Z}_{\geq 0}$ and $\mu\in{\rm PGL}_{2}(\mathbb{C})$ such that $X=F^{\circ m}\circ\mu$ and $B=\mu^{-1}\circ F^{\circ k}\circ\mu$.

### v2

@@ 3.1. Pre-simple polynomials and the case $d\geq 4$
By Lemma 3.21 and Theorem 3.23, there exists a non-empty Zariski open subset $W_{d}$ of $\mathrm{Poly}^{d}(\mathbb{C})$ such that for every $f\in W_{d}$, the polynomial $f$ is pre-simple and
(3.5) $\{g\in\mathrm{Rat}_{d}(\mathbb{C})\colon\mu_{g}=\mu_{f}\}=\{f\}.$
We show that $W_{d}$ also satisfies (2). It suffices to prove the “only if” direction. Let $f\in W_{d}$ and $g\in\mathrm{Poly}^{d}(\mathbb{C})$ be a pre-simple polynomial intertwined with $f$. By Theorem 3.28, there exist $k\in\mathbb{Z}_{>0}$ and $\alpha\in\mathrm{Aff}(\mathbb{C})$ such that
@@
By (3.5), $f=\alpha\circ g\circ\alpha^{-1}$. The polynomials $f$ and $g$ are conjugate as rational maps, so they are affinely conjugate, i.e., $[f]=[g]$ in $\mathrm{MPoly}^{d}(\mathbb{C})$.
@@ 3.2. The case of degree $2\leq d\leq 3$
Having analyzed general polynomials of degree $d\geq 4$ in § 3.1 via the notion of pre-simplicity, we now turn to general polynomials of degree $2\leq d\leq 3$, a study largely motivated by Pakovich’s work [Pak25b] on general rational maps of degree $2\leq d\leq 3$. We show that for $d\in\{2,3\}$, there exists a non-empty Zariski open subset $U$ of $\mathrm{Poly}^{d}(\mathbb{C})$ such that for all $(f,g)\in U\times U$, the polynomials $f$ and $g$ are intertwined if and only if they are (affinely) conjugate. This result is also a direct corollary of [FG22, Theorems 3.51 and 3.52] (see [GNY19, Theorem 1.4]).
For $d=2$, the moduli space $\mathrm{MPoly}^{2}\cong\mathbb{A}^{1}$ has dimension $1$, and every $f\in\mathrm{Poly}^{2}(\mathbb{C})$ is conjugate to a unique polynomial of the form $\phi_{c}(z)=z^{2}+c$ with $c\in\mathbb{C}$. Thus, polynomials of degree $2$ are easier to study due to the simple normal form $\phi_{c}$. For $d=3$, the moduli space (and the parameter space) is also relatively simple. Note that every polynomial of degree $2\leq d\leq 3$ is indecomposable because $d$ is prime.
Let $f\in\mathbb{C}[z]\setminus\mathbb{C}$ be a polynomial of degree $d\geq 2$. Recall the definitions:
$\Sigma(f):=\{\sigma\in\mathrm{PGL}_{2}(\mathbb{C})\colon f\circ\sigma=f\}\quad\text{and}\quad\Sigma_{\infty}(f):=\bigcup_{k\geq 1}\Sigma(f^{\circ k}).$
Compared with the case $d\geq 3$ (see Lemma 3.6), the case $d=2$ is significantly different:
@@
For every $c\in\mathbb{C}^{\times}$, the polynomial $\phi_{c}(z)=z^{2}+c$ satisfies $\Sigma(\phi_{c})=\Sigma_{\infty}(\phi_{c})=\{z,-z\}$, which is the cyclic group of order two generated by $z\mapsto-z$.
@@
Fix $c\neq 0$. Observe that $\phi_{c}$ is not conjugate to $z^{2}$.
Clearly, $\{z,-z\}\subseteq\Sigma(\phi_{c})\subseteq\Sigma_{\infty}(\phi_{c})$. Let $\sigma\in\Sigma(\phi_{c})$, and write $\sigma(z)=az+b$ with $a\in\mathbb{C}^{\times}$ and $b\in\mathbb{C}$ by (3.2). The equality $\phi_{c}\circ\sigma=\phi_{c}$ implies $\sigma(z)=\pm z$. Hence, $\Sigma(\phi_{c})=\{z,-z\}$.
We now show by induction on $n\in\mathbb{Z}_{>0}$ that $\Sigma(\phi_{c})=\Sigma(\phi_{c}^{\circ n})$ for all $n\geq 1$, which implies $\Sigma(\phi_{c})=\Sigma_{\infty}(\phi_{c})$. The case $n=1$ is trivial. Assume $\Sigma(\phi_{c})=\Sigma(\phi_{c}^{\circ(n-1)})$ for some $n\geq 2$. Clearly, $\Sigma(\phi_{c})\subseteq\Sigma(\phi_{c}^{\circ n})$. Let $\sigma\in\Sigma(\phi_{c}^{\circ n})$, and write $\sigma(z)=az+b$ with $a\in\mathbb{C}^{\times}$ and $b\in\mathbb{C}$ by (3.2). By [Pak25b, Lemma 3.4(ii)] and the identity
$\phi_{c}\circ\phi_{c}^{\circ(n-1)}=\phi_{c}\circ(\phi_{c}^{\circ(n-1)}\circ\sigma),$
either $\phi_{c}^{\circ(n-1)}=\phi_{c}^{\circ(n-1)}\circ\sigma$ or $\phi_{c}^{\circ(n-1)}=-\phi_{c}^{\circ(n-1)}\circ\sigma$.
If $\phi_{c}^{\circ(n-1)}=\phi_{c}^{\circ(n-1)}\circ\sigma$, then $\sigma\in\Sigma(\phi_{c}^{\circ(n-1)})=\Sigma(\phi_{c})$, so $\Sigma(\phi_{c}^{\circ n})\subseteq\Sigma(\phi_{c})$.
Suppose $\phi_{c}^{\circ(n-1)}=-\phi_{c}^{\circ(n-1)}\circ\sigma$. Comparing the coefficients of $z^{2^{n-1}}$, $z^{2^{n-1}-1}$, and $z^{2^{n-1}-2}$ yields:
  • $a^{2^{n-1}}=-1$;
  • $b=0$ (since the term $z^{2^{n-1}-1}$ vanishes in the expansion of $\phi_{c}^{\circ(n-1)}$);
  • $2^{n-2}c=-a^{2^{n-1}-2}2^{n-2}c$.
Since $c\neq 0$, the third equality implies $a^{2^{n-1}-2}=-1$. Then
@@
which is a contradiction.
Hence, $\Sigma(\phi_{c}^{\circ n})=\Sigma(\phi_{c})$. By induction, we obtain $\Sigma(\phi_{c})=\Sigma_{\infty}(\phi_{c})$. ∎
@@
There exists a non-empty Zariski open subset $U$ of $\mathrm{Poly}^{2}(\mathbb{C})$ such that for every $F\in U$:
@@
    $\Sigma(F)=\Sigma_{\infty}(F)$ is cyclic of order two;
@@
    if $G(z)\in\mathbb{C}(z)$ is a rational map of degree $2$ such that $G^{\circ k}=F^{\circ k}$ for some $k\in\mathbb{Z}_{>0}$, then $G=F$.
@@
Let $U:=\{a_{1}^{2}-2a_{1}\neq 4a_{0}a_{2}\}$, which is a non-empty Zariski open subset of
$\mathrm{Poly}^{2}(\mathbb{C})=\{a_{0}z^{2}+a_{1}z+a_{2}\}\cong\{(a_{0},a_{1},a_{2})\in\mathbb{C}^{3}\colon a_{0}\neq 0\}.$
A simple computation shows that $a_{0}z^{2}+a_{1}z+a_{2}\in\mathrm{Poly}^{2}(\mathbb{C})$ is conjugate to some $z^{2}+c$ with $c\in\mathbb{C}^{\times}$ if and only if its coefficients lie in $U$.
Let $F(z)=a_{0}z^{2}+a_{1}z+a_{2}\in U$, and let $c=c(F)$ be the unique number in $\mathbb{C}^{\times}$ such that $F$ is conjugate to $\phi_{c}$. Choose $\tau\in\mathrm{Aff}(\mathbb{C})$ with $\tau\circ F\circ\tau^{-1}=\phi_{c}$.
(1) By Lemma 3.30, $\Sigma(\phi_{c})=\Sigma_{\infty}(\phi_{c})=\{z,-z\}$ is cyclic of order two, so $\Sigma(F)=\Sigma_{\infty}(F)$ is also cyclic of order two.
(2) Let $G\in\mathbb{C}(z)$ be a rational map of degree $2$ with $G^{\circ k}=F^{\circ k}$ for some $k\in\mathbb{Z}_{>0}$. By the construction of $U$ and Lemma 3.13, $G\in C(F^{\circ k})$ must be a polynomial (of degree $2$). Set $\tilde{G}=\tau\circ G\circ\tau^{-1}\in\mathbb{C}[z]$. Then $\tilde{G}^{\circ k}=\phi_{c}^{\circ k}$, and we conclude that $\tilde{G}$ (and $G$) is not conjugate to $z^{2}$. By (1) and Lemma 3.30,
$\Sigma(\tilde{G})=\Sigma_{\infty}(\tilde{G})=\Sigma(\tilde{G}^{\circ k})=\Sigma(\phi_{c}^{\circ k})=\Sigma(\phi_{c})=\{\pm z\}.$
By [Pak25b, Lemma 3.5], $\tilde{G}=\nu\circ\phi_{c}$ for some $\nu\in\mathrm{PGL}_{2}(\mathbb{C})$. In fact, $\nu\in\mathrm{Aff}(\mathbb{C})$ because $\tilde{G}$ is a polynomial. From $\phi_{c}^{\circ k}=\tilde{G}^{\circ k}=(\nu\circ\phi_{c})^{\circ k}$, we get $\phi_{c}^{\circ(k-1)}=(\nu\circ\phi_{c})^{\circ(k-1)}\circ\nu$, so
@@
The same argument shows that $\Sigma(\phi_{c}\circ\nu)=\{\pm z\}$. On the other hand, by (3.1), we have $\Sigma(\phi_{c}\circ\nu)=\nu^{-1}\circ\Sigma(\phi_{c})\circ\nu=\{z,\nu^{-1}(-\nu(z))\}$. So $\nu^{-1}(-\nu(z))=-z$, which implies $\nu(-z)=-\nu(z)$. Thus, $\nu(z)=\lambda z$ for some $\lambda\in\mathbb{C}^{\times}$. Then
@@
Comparing the coefficient of $z^{2^{k}}$ on the two sides of (3.6) gives $\lambda^{2^{k}-1}=1$. If $k=1$, then $\lambda=1$. If $k\geq 2$, the coefficient of $z^{2^{k}-4}$ gives
@@
which implies $\lambda^{2}=1$ (since $c\neq 0$, $\lambda^{2^{k}-1}=1$, and $\lambda^{2^{k}-3}$ must equal 1). Then $\lambda=(\lambda^{2})^{2^{k-1}}/\lambda^{2^{k}-1}=1$. In all cases, we obtain $\nu(z)=z$, so $\tilde{G}=\phi_{c}$ and $G=F$. ∎
@@
There exists a non-empty Zariski open subset $U$ of $\mathrm{Poly}^{3}(\mathbb{C})$ such that for every $F\in U$:
@@
    if $G(z)\in\mathbb{C}(z)$ is a rational map of degree $3$ such that $G^{\circ k}=F^{\circ k}$ for some $k\in\mathbb{Z}_{>0}$, then $G=F$.
@@
By Lemmas 3.21 and 3.22, there exists a non-empty Zariski open subset $U$ of $\mathrm{Poly}^{3}(\mathbb{C})$ such that every $F\in U$ is pre-simple and $G(F)=1$. (In particular, every $F\in U$ is not conjugate to $z^{3}$.) Let $F\in U$. We prove (1) and (2).
(1) $F$ is pre-simple by assumption. Then $\Sigma(F)=\Sigma_{\infty}(F)=1$ by Corollary 3.9.
(2) Let $G\in\mathbb{C}(z)$ be a rational map of degree $3$ with $G^{\circ k}=F^{\circ k}$ for some $k\in\mathbb{Z}_{>0}$. The case $k=1$ is trivial. We assume $k\geq 2$. By Lemma 3.13, $G\in C(F^{\circ k})$ must be a polynomial. Applying Proposition 3.11 to the complete decomposition $F^{\circ k}=G\circ\cdots\circ G$, we conclude that
$G=\nu^{-1}\circ F=\delta^{-1}\circ F\circ\nu$
for some $\nu,\delta\in\mathrm{Aff}(\mathbb{C})$. (Here $\delta(z)=z$ when $k=2$.) Then $F\circ\nu=(\delta\circ\nu^{-1})\circ F$, so $\nu\in G(F)=1$. We conclude that $\nu(z)=z$ and $G=\nu^{-1}\circ F=F$. ∎
@@
For $d\geq 4$, the result follows from Lemmas 3.24 and 3.21.
If $d=2$, then $F$ is a generalized Lattès map if and only if it is exceptional (see Proposition 2.4). Exceptional polynomials of degree $2$ are exactly those polynomials conjugate to $z^{2}$ or $T_{2}(z)=z^{2}-2$. Since $\mathrm{MPoly}^{2}(\mathbb{C})\cong\mathbb{A}^{1}(\mathbb{C})$ has dimension $1>0$, a general $F$ is not exceptional.
If $d=3$, then $F$ is exceptional if and only if it is (affinely) conjugate to $z^{3}$ or $\pm T_{3}(z)=\pm(z^{3}-3z)$. Thus, a general $F$ is not exceptional. By Proposition 2.4, a non-exceptional $F$ is a generalized Lattès map if and only if it is (affinely) conjugate to
$zR_{a,b}(z)^{2}=z(az+b)^{2}$
for some degree one polynomial $R_{a,b}(z)=az+b$ with $a,b\in\mathbb{C}^{\times}$. Thus, $F$ is a generalized Lattès map if and only if there exist $a,b,e\in\mathbb{C}^{\times}$ and $f\in\mathbb{C}$ such that
(3.7) $F(z)=(ez+f)^{-1}\circ R_{a,b}(z)\circ(ez+f)=\frac{(ez+f)(a(ez+f)+b)^{2}-f}{e}.$
@@
There are only $3$ parameters $\lambda,b\in\mathbb{C}^{\times}$ and $\mu\in\mathbb{C}$, while
$\mathrm{Poly}^{3}(\mathbb{C})=\{a_{0}z^{3}+a_{1}z^{2}+a_{2}z+a_{3}\colon(a_{0},a_{1},a_{2},a_{3})\in\mathbb{C}^{\times}\times\mathbb{C}^{3}\}$
has dimension $4>3$. Hence, a general $F$ is not a generalized Lattès map. ∎
@@
Let $d\in\{2,3\}$. There exists a non-empty Zariski open subset $U$ of $\mathrm{Poly}^{d}(\mathbb{C})$ such that for every $F\in U$, if $B,X\in\mathbb{C}(z)\setminus\mathbb{C}$ and $k\in\mathbb{Z}_{>0}$ satisfy $F^{\circ k}\circ X=X\circ B$, then there exist $m\in\mathbb{Z}_{\geq 0}$ and $\mu\in\mathrm{PGL}_{2}(\mathbb{C})$ such that $X=F^{\circ m}\circ\mu$ and $B=\mu^{-1}\circ F^{\circ k}\circ\mu$.